An aerosol can is at a pressure of 8.7 atm at a temperatue of 88.0 C. It is then placed in a fire and the pressure rises to 20.8 atm. What is the temperature of the fire? Express your answer in degrees celcius, C.


1
Expert's answer
2021-05-13T04:57:46-0400

Using Gay-Lussac’s Law ;"\frac {P_1} {T_1} =\frac {P_2} {T_2}"


"T_2=\frac {361 ×20.8}{8.7}=863.1"


"T_2=590\u00b0C"